A normal force 35N act over an area of 5m2. What is the pressure exerted by the force?

a. 7.00 dynes b. 7.00 joules c. 7.00 pascals d. 7.00 watts



1
Expert's answer
2021-04-19T17:09:20-0400

By defintion, the pressure is:


"p = \dfrac{F}{A}"

where "F = 35N" is the force, and "A = 5m^2" is the area over which the force acts. The units of the pressure are pascals. Thus, obtain:


"p = \dfrac{35N}{5m^2} = 7Pa"

Answer. c. 7.00 pascals.